# On the Quest for Effectiveness in Human Oversight: Interdisciplinary Perspectives

### v1

@@ 7. Conclusion
This paper has proposed an approach to effective human oversight – encompassing causal power, epistemic access, self-control, and fitting intentions – to remedy the lack of clarity surrounding the concept of effectiveness. This is a joint contribution of researchers in psychology, law, philosophy, and computer science. We argued that, in essence, a morally responsible overseer with fitting intentions is generally suitable for mitigating risks associated with high-risk AI systems. We have identified facilitators and inhibitors of effectiveness in three categories, namely the technical design of the system, individual characteristics of human overseers, and the environmental circumstances in which oversight occurs. Thereby, we have provided inspiration for steps towards a successful implementation of effective human oversight. We also discussed the extent to which our understanding of effective human oversight aligns with the European Union’s upcoming AI Act, in particular Article 14 on human oversight. Our analysis suggests that the AI Act could benefit from incorporating more nuanced and structured conditions, such as those proposed in this paper.
In closing, we want to advocate for a balanced approach to effective human oversight. Without a concrete idea on what effectiveness in human oversight is, no productive discussion of it is possible. On the one hand, we urge to not throw out the baby with the bathwater by per se dismissing human oversight as a safeguard without the necessary conceptual clarity. On the other hand, caution is warranted as well when emphatically including a notion of effective human oversight in regulations and guidelines without this conceptual clarity. Our interdisciplinary approach is intended to aid both the advocates and the critics of human oversight. For many relevant oversight contexts, the four conditions of effectiveness seem to provide a better picture on when and how human oversight is useful. We, therefore, believe that these conditions could form a pivot for future research on human oversight.
Acknowledgements

### v2

@@ 7. Conclusion
This paper has proposed an approach to effective human oversight – encompassing causal power, epistemic access, self-control, and fitting intentions – to remedy the lack of clarity surrounding the concept of effectiveness. This is a joint contribution of researchers in psychology, law, philosophy, and computer science. We argued that, in essence, a morally responsible oversight person with fitting intentions is generally suitable for mitigating risks associated with high-risk AI systems. We have identified facilitators and inhibitors of effectiveness in three categories, namely the technical design of the system, individual characteristics of oversight persons, and the environmental circumstances in which oversight occurs. Thereby, we have provided inspiration for steps towards a successful implementation of effective human oversight. We also discussed the extent to which our understanding of effective human oversight aligns with the European Union’s upcoming AI Act, in particular Article 14 on human oversight. Our analysis suggests that the AI Act could benefit from incorporating more nuanced and structured conditions, such as those proposed in this paper.
In closing, we want to advocate for a balanced approach to effective human oversight. Without a concrete idea of what effectiveness is in the context of human oversight, no productive discussion of it is possible. On the one hand, we urge to not throw out the baby with the bathwater by per se dismissing human oversight as a safeguard without the necessary conceptual clarity. On the other hand, caution is warranted as well when emphatically including a notion of effective human oversight in regulations and guidelines without this conceptual clarity. Our interdisciplinary approach is intended to aid both advocates and critics of human oversight. For many relevant oversight contexts, the four conditions of effectiveness seem to provide a better picture of when and how human oversight is useful. We, therefore, believe that these conditions could form a pivot for future research on human oversight.
Acknowledgements.